OpenStudy (anonymous):

calculate the following limit: 2n/(n+12) as n->infinity

OpenStudy (zarkon):

2

OpenStudy (zarkon):

\[\frac{2n}{n+12}=\frac{2}{1+\frac{12}{n}}\rightarrow\frac{2}{1+0}=2\]